B =X Y 1 17 3 15 5 10 7 7 9 1 HW 1 For...

B =X Y
1 17
3 15
5 10
7 7
9 1

HW 1 For the date above

1-1. Construct an Analysis of Regression table, below (Below is the regression table I made)

Source of Variation

Degrees of Freedom, d.f.

Sums of Squares, SS

Mean Squares, MS

F, FR, MSR/MSE

Total

4

164

Regression

1

160

160

120.3

Residual

3

4

1.33

(I need a help with this) 1-2. Test B= 0 with a = 1%
Answer #1

Hypotheses are:

H_{0}:\beta_{1}=0,H_{a}:\beta_{1}\neq 0

Following table shows the calculations;

X Y X^2 Y^2 XY
1 17 1 289 17
3 15 9 225 45
5 10 25 100 50
7 7 49 49 49
8 1 64 1 8
Total 24 50 148 664 169

Sample size: n =5

Now,

S_{yy}=\sum y^{2}-\frac{\left (\sum y \right )^{2}}{n}=164

S_{xx}=\sum x^{2}-\frac{\left (\sum x \right )^{2}}{n}=32.8

S_{xy}=\sum xy-\frac{\left (\sum x \right )\left (\sum y \right )}{n}=-71

Slope of the regression equation is

b_{1}=\frac{S_{xy}}{S_{xx}}=-2.1646

and intercept of the equation will be

b_{0}=\frac{1}{n}(\sum y - b_{1} \sum x)=20.3902

Let us find SSE first :

SSE=\sum y^{2}-b_{0}\sum y-b_{1}\sum xy=10.31097561

So standard error of estimate will be

S_{e}=\sqrt{\frac{SSE}{n-2}}=1.853912584

The standard error of slope is

s_{b_{1}}=\frac{S_{e}}{\sqrt{S_{xx}}}=0.3237

T-statistics is

t=\frac{b_{1}-0}{s_{b_{1}}}=-6.687

Degree of freedom of test is df=n-2=5-2=3

The p-value using excel function "=TINV(0.01,3)" are +/- 5.841.

Since t does not lie between the critical values so we reject the null hypothesis.